# Modelling two-dimensional droplet rebound off deep fluid baths

### v2

@@ 5.2 Conclusion
The lubrication-mediated model that we have introduced has the capability of capturing the height and pressure within the air layer through the solution of an elliptic thin film problem over a dynamic boundary, which is novel at this level of detail within the literature within the context of this problem. We are able to solve for solid impactors and for deformable droplets in the small deformation limit. Our model is also restricted to small bath deformations and small viscous effects.
We have shown through an in-depth comparison for a single rebound case that the lubrication-mediated model is able to qualitatively match with the DNS system, obtaining similar behaviours and dynamics. A highlight is the comparison of air layer thicknesses shown in the previous section, which has good agreement with the morphology of air layers seen in droplet rebound studies (Tang et al., 2019) and DNS.
Across a range of parameters we have shown the lubrication-mediated model is able to capture behaviour of models based on the kinematic match idea, matching qualitatively with those 3D reduced methods (Galeano-Rios et al., 2021; Alventosa et al., 2023), yet providing additional insight into the air layer dynamics without increased computational effort. We have discussed the numerical stiffness observed in our system, and have proposed viscous damping methods to mitigate this issue. Further investigation is needed to improve the method and adapt it to regimes which permit larger deformation of the drop and liquid bath.
Finally, we have demonstrated the robustness of the lubrication-mediated model, and its applicability to repeated bouncing behaviour under sustained forcing as developed by Couder et al. (2005), with the future avenue of work open to investigate the coupling of the lubrication mediated model to mode thorough Faraday pilot wave studies, such as those found in Bush (2015), Galeano-Rios et al. (2017) and Durey et al. (2020).
[…]
R.C. gratefully acknowledges support from US-CBET/UK-EPSRC grant EP/W016036/1.
@@ Appendix A Physical Parameters
Table 1 lists the parameters used in the paper.
@@
Drop radius (silicone) $R_{0}$ ${8.4}\times 10^{-4}$ m
@@
Appendix B Viscosity consideration
@@
We chose to resolve this issue with an artificially enhanced viscosity during contact, specifically while $h\leq\varepsilon R_{0}$, and describe here the choice of viscosity and its effect on the results of a test deformable case. An in-depth stability analysis of the system and other solutions for this numerical issue is left for future work. Briefly, one could consider fully implicit numerical schemes (which pose some challenges due to the nonlinearity of the elliptic lubrication equation), split step schemes where the drop oscillations and lubrication equation are resolved implicitly or further asymptotic reduction of the droplet equations duting impact.
Figure 12: Trajectories for droplets with $R_{0}=0.2\times 10^{-3}$m, and $W_{0}=-0.2$ms-1, using a modified viscosity $\mu*=A\mu$, with $A$ ranging from $17$ to $500$. The solid black line corresponds to DNS results, and the dashed black line corresponding to the solid rebound of the same size and initial velocity.
We introduce a viscosity enhancement coefficient $A$, such that $\mu^{*}=A\mu$ during contact. We determined the smallest value of $A$ for which the entire parameter sweep of simulations would be stable, through over-damping the droplet and slowly lowering the value of $A$. Figure 12 demonstrates a range of values of $A$ for the single run case with $R_{0}=0.2\times 10^{-3}$m and $W_{0}=-0.2$ms-1. We present the solid lubrication mediated model with corresponding values, alongside DNS of the same initial conditions. The solid model can be interpreted as the limit $A\to\infty$. For this case, the choice $A=17$ was the lowest value that could be used reliably without introducing numerical instabilities. By contrast, large values of $A>400$ seemed to converge to the same result. Considering the full $W\!e$ and $Oh$ parameter regime within our study, the choice $A=150$ provided the smallest damping without instability.

### v3

@@ 6 Conclusion
The lubrication-mediated model that we have introduced has the capability of capturing the height and pressure within the air layer through the solution of an elliptic thin film problem over a dynamic boundary, which is novel in this context. We are able to solve for solid impactors and for deformable droplets in the small deformation limit. Our model is also restricted to small bath deformations and small viscous effects.
We have shown through an in-depth comparison for a single rebound case that the lubrication-mediated model is able to qualitatively match with the DNS system, obtaining similar behaviours and dynamics. A highlight is the comparison of air layer thicknesses shown in the previous section, which has good agreement with the morphology of air layers seen in droplet rebound experimental [40] and DNS studies.
Across a range of parameters we have uncovered that the lubrication-mediated model is able to capture behaviour of models based on the kinematic match approach, matching qualitatively with the respective 3D reduced methods [29, 33], yet providing additional insight into the air layer dynamics without increased computational effort. We have discussed the numerical stiffness observed in our system, and have proposed viscous damping methods to mitigate this issue. Further investigation is needed to improve the method and adapt it to regimes which permit larger deformation of the drop and liquid bath.
There are several avenues to extend this work, and the related 3D rigid-sphere axisymmetric computations in [44]. Two direct extensions are three-dimensional droplet impacts and two-dimensional non-symmetric impacts.
Finally, we have demonstrated the robustness of the lubrication-mediated model, and its applicability to repeated bouncing behaviour under sustained forcing as first discovered by Couder et al [13]. Future avenues of work would be to investigate the results of the lubrication mediated model in Faraday pilot wave studies such as for walkers where almost vertical impacts capture the dynamics [31], and in dynamics where the vertical motion is known to play an important role such as in corrals [49], tunnelling [19] and non-resonant behaviour [48].
@@ Appendix A Physical parameters
Table 1 lists the parameters used in the present study.
@@
Drop radius (silicone) $R_{0}$ $8.3\times 10^{-4}$ m
@@
Appendix B Boundary layer scaling
Solutions to the quasi-potential model laid out by equations (6)-(14) exhibit a boundary layer effect at higher Reynolds numbers. In order to see that one may solve the (linear) ‘free-wave’ problem and obtain explicitly the forms of the free surface $\eta_{b}$ and velocity potential $\phi_{b}$ as
$\eta_{b}=ae^{i(kx\pm\omega_{0}t)}e^{-2\nu_{b}k^{2}t},\qquad\phi_{b}=\pm ia\frac{\omega_{0}}{|k|}e^{i(kx\pm\omega_{0}t)}e^{-2\nu_{b}k^{2}t}e^{|k|z},$
where $\omega_{0}=\sqrt{|k|(g+\sigma_{b}k^{2}/\rho_{b})}$ is the frequency of undamped capillary-gravity waves. As expected, this is a time decaying travelling wave. Furthermore, one may now compute the induced boundary layer represented by $\psi(x,z,t)$ using $\partial_{x}\psi(x,0,t)=2\nu_{b}\partial_{x}^{2}\eta_{b}(x,t)$ (from the leading order balance of (2.17)). This results in
$\Psi=2ik\nu_{b}ae^{i(kx\pm\omega_{0}t)}e^{-2\nu_{b}k^{2}t}e^{\pm i\sqrt{2}\sqrt{Re_{\omega}}|k|z/2}e^{\sqrt{2}\sqrt{Re_{\omega}}|k|z/2},$
where $Re_{\omega}=\frac{\omega_{0}}{\nu_{b}k^{2}}$. Hence one finds that the vertical decay length scale for $\psi$ is $O(Re_{\omega}^{-1/2})$ and that denoting $u^{\psi},v^{\psi}$ and $u^{\phi},v^{\phi}$ as the vortical and irrotational components of the velocity field respectively, we have
$u^{\psi}\sim Re_{\omega}^{-1/2}u^{\phi},\qquad v^{\psi}\sim Re_{\omega}^{-1}v^{\phi}.$
This describes the boundary layer structure as $Re_{\omega}\rightarrow\infty$.
Appendix C Viscosity consideration
@@
We chose to resolve this issue with an artificially enhanced viscosity during contact, specifically while $h\leq\varepsilon R_{0}$, and describe here the choice of viscosity and its effect on the results of a test deformable case. An in-depth stability analysis of the system and other solutions for this numerical issue is left for future work. Briefly, one could consider fully implicit numerical schemes (which pose some challenges due to the nonlinearity of the elliptic lubrication equation), split step schemes where the drop oscillations and lubrication equation are resolved implicitly or further asymptotic reduction of the droplet equations during impact.
Figure 14: Trajectories for droplets with $R_{0}=0.2$mm, and $W_{0}=-0.2$ms-1, using a modified viscosity $\mu^{*}=A\mu$, with $A$ ranging from $17$ to $500$. The solid black line corresponds to DNS results, and the dashed black line corresponding to the solid rebound of the same size and initial velocity.
We introduce a viscosity enhancement coefficient $A$, such that $\mu^{*}=A\mu$ during contact. We determined the smallest value of $A$ for which the entire parameter sweep of simulations would be stable, through over-damping the droplet and slowly lowering the value of $A$. Figure 14 demonstrates a range of values of $A$ for the single run case with $R_{0}=0.2$mm and $W_{0}=-0.2$ms-1. We present the solid lubrication mediated model with corresponding values, alongside DNS of the same initial conditions. The solid model can be interpreted as the limit $A\to\infty$. For this case, the choice $A=17$ was the lowest value that could be used reliably without introducing numerical instabilities. By contrast, large values of $A>400$ seemed to converge to the same result. Considering the full $W\!e$ and $Oh$ parameter regime within our study, the choice $A=150$ provided the smallest damping without instability.
@@
Funding:
K.A.P is supported by a scholarship from the EPSRC Centre for Doctoral Training in Statistical Applied Mathematics at Bath (SAMBa), under the project EP/S022945/1. R.C. gratefully acknowledges support from US-CBET/UK-EPSRC grant EP/W016036/1.